a large pipe can fill a tank 10 min faster than it takes a smaller pipe to fill the same tank. together, both pipes can fill the tank in 12 min. how long would it take the large pipe working alone to fill the tank?
Answer by jorel1380(3719)

1/x+1/x-10=1/12
12(x-10)+12x=x2-10x
12x-120+12x=x2-10x
0=x2-34x+120
0=(x-30)(x-4)
x=30 or 4
Throwing out 4 as an answer, we get the small pipe taking 30 min., the large pipe taking 20 min..